x3y3(2ydx+ xdy) = (5ydx+7xdy)


1
Expert's answer
2021-02-24T12:47:10-0500

Solution.

"x^3y^3(2ydx+xdy)=5ydx+7xdy."

Divide both parts of the equation by "dx:"

"x^4y^3\frac{dy}{dx}+2x^3y^4-5y-7x\frac{dy}{dx}=0,"

or "x^4y^3y'+2x^3y^4-5y-7xy'=0."

Make a replacement

"u(x)=xy(x)," from here "y(x)=\frac{u(x)}{x}," and "y'=\frac{u'(x)}{x}-\frac{u(x)}{x^2}."

Substitute the obtained expressions in our equation:

"x^4\frac{u^3}{x^3}(\frac{u'}{x}-\frac{u}{x^2})+2x^3\frac{u^4}{x^4}-5\frac{u}{x}-7x(\frac{u'}{x}-\frac{u}{x^2})=0,"

or "u^3u'-7u'+\frac{u^4}{x}+\frac{2u}{x}=0."

Divide both parts of the equation by "u^4+2u" and write the equation in the form:


"\frac{(u^3-7)u'}{(u^3+2)u}=-\frac{1}{x}."

Thus we obtain a differential equation with dividing terms. Let's solve it:


"\frac{dx(u^3-7)\frac{du}{dx}}{(u^3+2)u}=-\frac{dx}{x},""\frac{du(u^3-7)}{(u^3+2)u}=-\frac{dx}{x},""\int\frac{u^3-7}{(u^3+2)u}du=\int(-\frac{1}{x})dx,
\newline
\int\frac{9u^2}{2(u^3+2)}du-\int\frac{7}{2u}du=\int(-\frac{1}{x})dx,""\frac{3}{2}\ln{|u^3+2|}-\frac{7}{2}\ln(|u|)=-\ln{|x|}+C,"

where C is some constant.

Returning to the replacement we obtain the solution of the given equation:


"\frac{3}{2}\ln{|x^3y^3(x)+2|}-\frac{7}{2}\ln(|xy(x)|)=-\ln{|x|}+C."

Answer.

"\frac{3}{2}\ln{|x^3y^3(x)+2|}-\frac{7}{2}\ln(|xy(x)|)=-\ln{|x|}+C."